# Know Unreported Roadway Incidents in Real-time: A Deep Learning Framework for Early Traffic Anomaly Detection

### v1

@@ 5.2.1 Detection Examples
Using I-70 as an example, our prediction process during the evening period on 2023-01-11 is detailed in Appendix A. In this example, we use our model trained with the ahead labelling strategy. In Appendix A, the red dashed line at 0.56 represents the adaptive threshold the validation set generated. At 17:15, our model’s predictions for the next six time steps are all below the threshold, so no alert is triggered (the blue line is used to indicate no alert). However, at 17:25, the minimum value exceeds the threshold, triggering an alert (the red line is used to indicate an alert). This alert ends at 17:45, as the minimum predicted value falls below the threshold at that time. The black text in Appendix A represents the current speed at the predicted time, while the red shading indicates the report time. From this, it can be observed that our model triggered an alert earlier than the incident report and even before the speed significantly dropped.
@@ 5.2.2 Examples of detecting anomalies and possibly unreported incidents
The above are two examples of direct prediction processes. During the periods mentioned, there were incident reports, and our model also issued alerts. However, there were also instances where no reports were present, but the alert was triggered (’False Positive’ Cases), and cases where an incident report appeared later, but our model did not trigger an alert (’False Negative’ Cases). Figures 13 and 14 illustrate these two kinds of cases, respectively.
@@ 5.2.3 Overall Results
Considering the issues caused by incident missing report, as shown in Figures 3 and 13, we have optimized the definition of FAR (False Alarm Rate) in Equation 13. The numerator of FAR has been changed from the number of alarms not coinciding with any incident reports to the number of alarms not coinciding with any incident reports or anomalies. The anomalies here are those generated using the anomaly labels described in Section 4.2. This adjustment reduces the likelihood of alarms that captured anomalies, such as those in Figure 13, being classified as false alarms.
@@
We adopted the following conventional AID baselines. It is important to note that, as we have already demonstrated the effectiveness of training with anomaly labels in the previous chapter, and since almost all existing studies use incident reports for training, all the baselines here, except for the first one whose detection process is based on statistical methods, are trained using incident reports. Given the differences in input data, for example, we do not have loop detector data, we cannot directly use the original baseline models. However, we have made efforts to retain the proposed model’s structure and adapt it by tuning it to our dataset.
@@
  • SVM, Support Vector Machine. We train a binary classification SVM with an RBF kernel based on whether there is an incident report in the next time step.
  • GAN, Generative Adversarial Network. We first use a GAN (Generative Adversarial Network) to learn the characteristics of samples with incident reports and generate some fake samples. These fake samples, along with the real samples, are used to train the SVM.
  • Ours. Our model trained without ahead-labelled anomaly labels
  • Ours (a). Our model trained with 15-minute ahead-labelled anomaly labels
In Table 4, we first crossed out some results that lack practical significance, such as extremely high FAR or cases where both DR and DR(S) are 0. The reasons for crossed out are marked in red.
@@
Among baseline methods, OOD only shows some practical significance in Howard County US-40, while other results face very high FAR and low DR. Notably, they all converge well in the speed prediction model. This highlights the drawbacks of statistical methods, where outliers in SND are not necessarily anomalies but could be due to model prediction errors and probe vehicle speed observation errors. Although statistical methods have performed well in some studies, probe vehicle data may still be too noisy to support this method. Regarding SVM, the results of the SVM are entirely lacking in practical significance. This may be due to the difficulty in learning from incident reports. Because when we trained using anomaly labels (See Tables 2 and 3), they achieved convergence in most cases.
@@
GAN achieved good convergence across all cases in Howard County. This may be because GAN learned the anomaly characteristics from the incident reports, and its generator produced samples similar to those we generated using Algorithm 1. By using incident reports directly for training, the DR of GAN in I-70E and I-695A was higher than our method, although their FAR was also relatively high. Overall, GAN’s MTTD was still worse than our method, reflecting the lack of early features in the incident reports. In addition, GAN performed poorly in Cranberry Township, even worse than SVM in I-79S and I-79N, indicating that the samples generated by GAN actually added noise. This suggests that GAN still struggles to learn incident characteristics in the Cranberry Map. In the future, we might consider using prior knowledge, such as slowdown speed, to guide the learning process of GAN.
@@ 6 Conclusion and Future Work
In this study, we address an early anomaly detection or prediction problem that is more practically useful and scalable than conventional AID models. We propose a method that combines incident reports and traffic domain knowledge, i.e., the correlation between slowdown speed and incident, to generate anomaly labels and train a deep learning model for early anomaly detection. This method aims to overcome inherent issues in incident reports, including false reports, insignificant incident reports, delayed reports, and missing reports. We adopted ahead labeling and multi-step prediction strategies, allowing the model to predict incidents earlier than traditional methods. Additionally, we applied various machine learning strategies to address the issues of insufficient anomaly samples and data imbalance. Unlike existing research in the field of AID, our method aims to capture a broader range of anomalies beyond after the fact incident reports (including early incident characteristics and unreported anomalies) and more importantly, offer early alerts of anomalies that potentially influence the traffic flow in the short term. Furthermore, our method is highly scalable as it relies solely on ubiquitously available data. We validated our model on ten road sections in two road networks, achieving results that surpassed the baselines. This experiment shows a good trade-off between precision and recall, and resulting anomaly alerts 4-42 min earlier than Waze reports.
@@
Although the data used in this study is theoretically available in real-time, we need to coordinate with data providers to ensure we can access the real-time data required by the model as early as possible. This will maximize the model’s advantage in early prediction/detection for potential future applications. Besides, this study also has certain limitations that can be improved. First, regarding the selection of prior knowledge, we mainly chose edges with a high correlation between the occurrence of incidents and high slowdown speeds. In fact, due to the topological structure, not all edges have this correlation. Bases on our definition of slowdown speed, this high correlation requires that the segment be relatively long and that most incidents do not occur at the very beginning of each segment. For those that do not exhibit this high correlation, we may need to consider redefining the slowdown speed or using other values, such as the travel time index. Semi-supervised learning is also a potential solution for label denoising. Moreover, our algorithm is also difficult to handle for segments where only one or two incidents occur per year, as the validation and test sets may not contain any incidents. This might require using simulator data to generate sufficient training samples and finetuning with real-world incident samples. Additionally, we can tune a reasonable ahead-labeling period, which is also a direction worth exploring in future research.

### v2

@@ 5.3.1 Detection Examples
Using I-70 as an example, our prediction process during the evening period on 2023-01-11 is detailed in Appendix A. In this example, we use the Transformer model trained with labels generated by Algorithms 1 and 2 and optimization method detailed in section 4.5 and 4.6. In Appendix A, the red dashed line at 0.56 represents the adaptive threshold the threshold-tuning set generated. At 17:15, our model’s predictions for the next six time steps are all below the threshold, so no alert is triggered (the blue line is used to indicate no alert). However, at 17:25, the minimum value exceeds the threshold, triggering an alert (the red line is used to indicate an alert). This alert ends at 17:45, as the minimum predicted value falls below the threshold at that time. The black text in Appendix A represents the current speed at the predicted time, while the red shading indicates the report time. From this, it can be observed that our model triggered an alert earlier than the incident report and even before the speed significantly dropped.
@@ 5.3.2 Examples of detecting anomalies and possibly unreported incidents
The above are two examples of direct prediction processes. During the periods mentioned, there were incident reports, and our model also issued alerts. However, there were also instances where no reports were present, but the alert was triggered (“False Positive” Cases), and cases where an incident report appeared later, but our model did not trigger an alert (“False Negative” Cases). Figures 16 and 17 illustrate these two kinds of cases, respectively.
@@ 5.3.3 Overall Results
Considering the issues caused by incident missing report, as shown in Figures 3 and 16, we have optimized the definition of FAR (False Alarm Rate) in Equation 13. The numerator of FAR has been changed from the number of alarms not coinciding with any incident reports to the number of alarms not coinciding with any incident reports or anomalies. The anomalies here are those generated using the anomaly labels described in Section 4.2. This adjustment reduces the likelihood of alarms that captured anomalies, such as those in Figure 16, being classified as false alarms.
@@
According to Section 2, we adopted the following conventional AID baselines, including statistical, comparative, and AI-based methods. It is important to note that, as we have already demonstrated the effectiveness of training with anomaly labels in the previous chapter, and since almost all existing studies use incident reports for training, all the baselines here, except for the statistical (OOD) and comparative method (CFOD), are trained using incident reports. Given the differences in input data, for example, we do not have loop detector data, we cannot directly use the original baseline models. However, we have made efforts to retain the proposed model’s structure and adapt it by tuning it to our dataset.
@@
  • CFOD, Chebyshev Distance First Order Directive. According to the early-detection method proposed by Grigorev et al. (2024), for a given road segment, we first compute the speed difference between its recent input speed sequence and its corresponding monthly profile (i.e., the same day of the week and the time of the day over the past month). We then calculate the Chebyshev distance based on this difference. After normalizing the distance using parameters derived from the validation set, we compute its first-order derivative. The incident is considered to begin at a significant positive peak in the derivative and end at a substantial negative peak.
  • SVM, Support Vector Machine. Inspired by the idea proposed by Yuan and Cheu (2003), we train a binary classification SVM with an RBF kernel based on whether there is an incident report in the next time step.
  • GAN, Generative Adversarial Network. Inspired by the idea proposed by Lin et al. (2020), we first use a GAN (Generative Adversarial Network) to learn the characteristics of samples with incident reports and generate some fake samples. These fake samples, along with the real samples, are used to train the SVM.
  • TSSAE, Temporal and Spatially Stacked Autoencoder. The model architecture is proposed by Li et al. (2022). To adapt to our data, we fed all relevant input features of each segment within the sub-graph into a separate autoencoder. The autoencoders were first pre-trained in an unsupervised manner and then fine-tuned using incident labels. For each segment, we evaluated two settings—with and without GAN-based data augmentation—and reported the best-performing results.
  • Ours. The proposed method using the Transformer model trained without ahead-labeled anomaly labels.
  • Ours (a). The proposed method using the Transformer model trained with 15-minute ahead-labeled anomaly labels.
In Table 4, we first crossed out some results that lack practical significance, such as extremely high FAR ($>$0.8) or cases where both DR and DR(S) are 0. The reasons for crossed out are marked in red.
@@
Among baseline methods, OOD only shows some practical significance in Howard County US-40, while other results face very high FAR and low DR. Notably, they all converge well in the speed prediction model. This highlights the drawbacks of statistical methods, where outliers in SND are not necessarily anomalies but could be due to model prediction errors and probe vehicle speed observation errors. Although statistical methods have performed well in some studies, probe vehicle data may still be too noisy to support this method.
Regarding CFOD, on the Howard County map, we observe that CFOD suffers from a significantly low detection rate and a relatively high false alarm rate (FAR), especially on I-70E, I-70W, and US-40W segments. This suggests that relying solely on temporal differences may be insufficient for effective incident detection. Furthermore, CFOD performs poorly on the Cranberry map, implying that incidents in Cranberry may exhibit less temporal fluctuation and instead manifest primarily through spatial patterns, as explained in section 4.2, which CFOD cannot capture. Although CFOD shows limited overall performance, it reveals an interesting insight: in the US-40W and I-695A segments of Howard County, CFOD is able to identify certain incident reports not labeled as significant in our dataset. This suggests that our labeling criteria may miss incidents with localized, segment-level temporal anomalies, which could be valuable for refining future labeling strategies. Moreover, given the strong theoretical interpretability of CFOD, we further analyzed its false alarms and found that most of them were caused by point anomalies—i.e., anomalies that occur at only a single time point. These point anomalies may result from noisy input data or may correspond to incidents that have limited impact on traffic conditions and were therefore not labeled as significant.
@@
Regarding SVM, the results of the SVM are entirely lacking in practical significance. This may be due to the difficulty in learning from incident reports. This inference is supported by our observation that, when trained using anomaly labels (see Tables 2 and 3), the models were able to converge in most cases—suggesting that the quality of incident report labels may hinder effective learning. GAN achieved good convergence across all cases in Howard County. This may be because GAN learned the anomaly characteristics from the incident reports, and its generator produced samples similar to those we generated using Algorithm 1. By using incident reports directly for training, the DR of GAN in I-70E and I-695A was higher than our method, although their FAR was also relatively high. Regarding MTTD, the GAN achieved the best performance on I-76. However, it is important to note that the GAN’s DR on this segment is very low, indicating that the low average MTTD may result from only a few detected incidents. Moreover, the FAR is high—approaching our unacceptable threshold of 0.8—which further limits the practical significance of these MTTD results. Overall, GAN’s MTTD was still worse than our method, reflecting the lack of early features in the incident reports. In addition, GAN performed poorly in Cranberry Township, even worse than SVM in I-79S and I-79N, indicating that the samples generated by GAN actually added noise. This suggests that GAN still struggles to learn incident characteristics in the Cranberry Map. In combination with the CFOD results, it is reasonable to infer that the GAN may struggle to learn spatial anomalies effectively. In the future, we might consider using prior knowledge, such as slowdown speed, or by designing a discriminator with a graph-based structure, to guide the learning process of GAN.
As for TSSAE, the results are similar to those of SVM in that the model fails to converge. While we observed clear convergence during the pretraining stage, the model showed little to no convergence during finetuning. We also attempted to replace the architecture with the Transformer, but it similarly failed to converge. This further suggests that under noisy labeling conditions, simply improving the model architecture is insufficient to enhance performance. Notably, as in the original paper, we also attempted GAN-based data augmentation prior to training. However, the model still failed to converge and performed worse than the GAN-SVM combination. This likely reflects limitations in both data volume and label quality, which are insufficient to support effective deep learning model training even fine-tuning.
5.4 Sensitivity Analysis on Slowdown Speed Threshold
$\theta_{\text{SD}}^{\text{I-70E}}$ (mph) Ahead-Labeling DR MTTD FAR DR(S) MTTD(S)
30 0.46 -7 0.00 1.00 -7
30 0.46 -12 0.23 1.00 -12
25 0.61 -7 0.15 1.00 -7
25 0.61 -9 0.20 1.00 -9
20 0.46 -6 0.11 0.75 -6
20 0.92 -17 0.34 1.00 -24
Table 5: Sensitivity Analysis
In Algorithm 1, we set $\theta_{1}$ and $\theta_{2}$ to determine the slowdown speed threshold $\theta_{\text{SD}}^{i}$. An incident report is marked as significant if it causes a slowdown speed value higher than this threshold. Additionally, if the slowdown speed continuously remains above $\theta_{\text{SD}}^{i}$, it is also considered anomalous. We now present how the choice of threshold impacts the model’s performance. It is important to note that since the threshold selection directly affects the ground truth labels for anomalies, direct comparisons using metrics such as those in Table 2 or Table 3 are not meaningful. Instead, similar to Table 4, we compare the detection results against incident reports. However, care must be taken when interpreting metrics like FAR, DR(S), and MTTD(S), as different thresholds result in different sets of significant anomalies or incident reports, which makes direct comparisons potentially misleading.
Take the I-70 segment as an example. As shown in Table 5, in the absence of ahead labeling, reducing $\theta_{\text{SD}}$ from 30 mph to 25 mph (i.e., comparing Row 1 and Row 3) introduces more incident reports into the training set. This leads to a notable increase in the Detection Rate (DR) but also causes a rise in the False Alarm Rate (FAR) since the inclusion of less severe incidents makes the model more likely to misclassify normal conditions as anomalies. Under the setting with ahead labeling, the DR also improves, while the FAR decreases numerically, which seems to be in contrast to the case with no ahead-labeling. Although the FAR does not appear to increase, this is mainly due to our definition, which states that the denominator of FAR is the total number of alerts. Lowering the threshold introduces more mild anomalies into the training set, making the model more sensitive and likely to trigger more alerts during inference. However, many of these alerts are considered anomalies under the relaxed threshold, thereby diluting the FAR.
When the slowdown speed threshold is further reduced to 20 mph, the model—without ahead labeling (Row 5)—fails to identify some incidents marked as significant (i.e., DR(S) is no longer 1). This is because the lower threshold introduces more anomalous samples that are difficult to distinguish from normal cases, even normal cases, into the training labels. This results in a drop in overall DR compared to the case where the threshold is set to 25 mph (Row 3). Compared to the threshold of 30 mph (Row 1), although the DR remains the same, the FAR is noticeably higher. When ahead-labeling is introduced, the proportion of incident samples in the training set increases, leading the model to make more aggressive predictions. However, this also results in more mild anomalies—or even normal samples—being mislabeled as anomalies, reducing the purity of training labels. Consequently, although DR increases significantly, the FAR also rises markedly.
Finally, MTTD does not show a significant change as the threshold decreases, which highlights the importance of ahead labeling for enabling early detection/prediction.
@@ 6 Conclusion and Future Work
In this study, we address an early anomaly detection or prediction problem that is more practically useful and scalable than conventional AID models. We propose a method that combines incident reports and traffic domain knowledge, i.e., the correlation between slowdown speed and incident, to generate anomaly labels and train a deep learning model for early anomaly detection. This method aims to overcome inherent issues in incident reports, including false reports, insignificant incident reports, delayed reports, and missing reports. We adopted ahead labeling and multi-step prediction strategies, allowing the model to predict incidents earlier than traditional methods. We also applied various machine learning strategies to address the issues of insufficient anomaly samples and data imbalance. Unlike existing research in the field of AID, our method aims to capture a broader range of anomalies beyond after the fact incident reports (including early incident characteristics and unreported anomalies) and, more importantly, it provides early alerts for anomalies that may disrupt traffic flow in the short term with regional impacts. Furthermore, our method is highly scalable as it relies solely on ubiquitously available data. We validated our model on ten road sections in two road networks, achieving results that surpassed the baselines. This experiment shows a good trade-off between precision and recall, resulting in anomaly alerts 5-42 min earlier than Waze reports.
@@
Although the data used in this study is theoretically available in real-time, we need to coordinate with data providers to ensure we can access the real-time data required by the model as early as possible. This will maximize the model’s advantage in early prediction/detection for potential future applications. Besides, this study also has certain limitations that can be improved. First, regarding the selection of prior knowledge, we mainly chose edges with a high correlation between the occurrence of incidents and high slowdown speeds. In fact, due to the topological structure, not all edges have this correlation. Bases on our definition of slowdown speed, this high correlation requires that the segment be relatively long and that most incidents do not occur at the very beginning of each segment. For those that do not exhibit this high correlation, we may need to consider redefining the slowdown speed or using other values, such as the travel time index. Although most case studies demonstrate that our method effectively captures both spatial and temporal anomalies—and clearly outperforms approaches that focus solely on temporal patterns—comparisons with baselines suggest that some anomalies with purely temporal, segment-level effects may still be overlooked. Future research could explore more integrated labeling strategies that jointly capture both spatial and temporal anomalies in a cohesive manner. Semi-supervised learning or agent-based learning will be a potential solution for this issue. Moreover, our algorithm is also difficult to handle for segments where only one or two incidents occur per year, as the validation and test sets may not contain any incidents. This might require using simulator data to generate sufficient training samples and finetuning with real-world incident samples. Additionally, we can tune a reasonable ahead-labeling period, which is also a direction worth exploring in future research.